OpenStudy (anonymous):

Solve the given equation. (Enter your answers as a comma-separated list. Let k be any integer. Round terms to three decimal places where appropriate. If there is no solution, enter NO SOLUTION.) 9 sin 2θ − 2 sin θ = 0

OpenStudy (anonymous):

Are there any restrictions? Such as a domain?

OpenStudy (anonymous):

no, answer will likely be (something) +2kpi

OpenStudy (anonymous):

9 sin (2theta) - 2 sin(theta) = 0 9 * 2 sin(theta)cos(theta) - 2 sin(theta) = 0 2sin(theta) * (9 cos(theta) - 1) = 0 2sin(theta) = 0 9cos(theta) - 1 = 0 sin(theta) = 0 cos(theta) = 1/9 sin = pi*k cos = cos^1(1/9) = +/- 1.459 + 2*k*pi answer: pi*k, 1.459 + 2*pi*k, -1.459 + 2*pi*k